OpenStudy (anonymous):

Of the 850 students at Brown Middle School, 38% are in the school band. How many students are in the school band? 32 527 323 812

OpenStudy (anonymous):

just take .38 times 850

OpenStudy (anonymous):

when doing percentages like this and the last one just take the decimal times the main number

jimthompson5910 (jim_thompson5910):

Part/Whole = Percent/100 x/850 = 38/100 x = (38/100)*850 x = (0.38)*850 x = ??? So it's following the same thing __assassin__ is saying, just in a slightly different way.

OpenStudy (anonymous):

thats --exactly-- what i said...

jimthompson5910 (jim_thompson5910):

not exactly, but along the same lines

OpenStudy (anonymous):

no it really is, you took the percentage, made it into a decimal, then multiplied by the same number

jimthompson5910 (jim_thompson5910):

no it's slightly different, yes it leads to the same thing it just has a different beginning to it